Sales at LL Boutique decreased 10% this month compared to last month. If sales this month were $130,581, what were the sales (in $) last month?


1
Expert's answer
2021-09-20T15:21:41-0400

Sales figure for last month = "x"

Sales figure for this month = $130581

Since sales dropped by 10% compared to last month, then

10% = 0.1


"\frac{x - 130581}{x} = 0.1"

"x-130581 = 0.1x"

"x-0.1x = 130581"

"0.9x = 130581"

= $145090

The sales figure for last month were $145090